solving systems of equations by substitution
1) y=-2x+2
y=1/3 x-5

Answer by ewatrrr(24785)

y=-2x+2 and y=1/3 x-5
Therefore
-2x+2 = (1/3)x-5
-6x + 6 = x - 15
21 = 7x
x = 3